I have submitted this question before and received an incorrect answer. Please help!

Triathlon times. In triathlons, it is common for racers to be placed into age and gender groups. Friends Leo and Mary both completed the Hermosa Beach Triathlon, where Leo competed in the Men, Ages 30 - 34 group while Mary competed in the Women, Ages 25 - 29 group. Leo completed the race in 4781 seconds, while Mary completed the race in 5591 seconds. Obviously Leo finished faster, but they are curious about how they did within their respective groups. Can you help them? Round all calculated answers to four decimal places.

Here is some information on the performance of their groups:

The finishing times of the Men, Ages 30 - 34 group has a mean of 4318 seconds with a standard deviation of 582 seconds.

The finishing times of the Women, Ages 25 - 29 group has a mean of 5240 seconds with a standard deviation of 771 seconds.

The distributions of finishing times for both groups are approximately Normal.

Remember: a better performance corresponds to a faster finish.

5. What percent of the triathletes did Leo finish slower than in his group?  %.

6. What percent of the triathletes did Mary finish faster than in her group?  %.

7. What is the cutoff time for the fastest 23% of athletes in the men's group, i.e. those who took the shortest 23% of time to finish?

seconds

8. What is the cutoff time for the slowest 32% of athletes in the women's group?

seconds

Incorrect answer:

5. 78.81

6. 32.38

7. 4748.098

8. 4879.172
Answer #1

5)

probability =P(X<4781)=(Z<4781-4318)/582)=P(Z<(0.7955)=0.7868~ 78.68 %

6)

probability =P(X>5591)=P(Z>(5591-5240)/771)=P(Z>0.46)=1-P(Z<0.46)=1-0.6755=0.3245~ 32.45 %

7)

for 23th percentile critical value of z= -0.739
therefore corresponding value=mean+z*std deviation= 3887.9911

8)

for 68th percentile critical value of z= 0.468
therefore corresponding value=mean+z*std deviation= 5600.5958